OpenStudy (anonymous):

how do you solve (x+2)^3/4=27?

OpenStudy (tkhunny):

I'm guessing "3/4" is the exponent? Please remember your Order of Operations and use parentheses where necessary to clarify intent. Have you considered raising both sides to the 4/3 power?

OpenStudy (anonymous):

\[\left( x+2 \right)^{\frac{ 3 }{ 4 }}=27\] \[\left[ \left( x+2 \right)^{\frac{ 1 }{ 4 }} \right]^{3}=3^{3}\] \[\left( x+2 \right)^{\frac{ 1 }{ 4 }}=3\] now raise power 4 on both sides and get the solution.

OpenStudy (anonymous):

Thanks!!